"The scouts sold small hoagies for $3 each and large hoagies for $4. if they sold a total of 260 hoagies worth $930 how many of each size hoagie did they sell
Answer by checkley71(8403)

S+L=260 S=260-L
3S+4L=930 SUBSTITUTING (260-L) FOR S IN THE SECOND EQUATION WE GET
3(260-L)+4L=930
780-3L+4L=930
L=930-780
L=150 SMALL HOAGIES SOLD
S=260-150
S=110 LARGE HOAGIES SOLD
PROOF
3*110+4*150=930
330+600=930
930=930